# What drives the growth of black holes: a decade of progress

### v1

@@ 5.4.1 Increasing the demographics of high-redshift AGN
Many LRDs have a spectral break around the wavelength of the Balmer break (Labbe et al., 2024; Ma et al., 2024; Wang et al., 2024). Analysing a large sample of LRD JWST spectra, Setton et al. (2024) found the inflexion in the “V” shaped SED corresponds to the Balmer break in 50% of sources; see Fig. 26 for the break in the composite LRD spectrum. They argued this break is unlikely to occur for dust-reddened AGN with a range of dust extinctions and host-galaxy strengths and, adopting “Occam’s razor”, proposed the properties of LRDs are from a single origin, potentially due to late-type stars. However, a growing number of studies have argued that an AGN buried in an extremely dense cocoon of gas (what is sometimes referred to as a “black-hole star” scenario) would also produce such feature and could simultaneously explain the Balmer absorption lines seen in the spectra of some JWST AGN (e.g., Inayoshi and Maiolino, 2024; de Graaff et al., 2025; Ji et al., 2025; Naidu et al., 2025). A potential consequence of the large gas density is scattering of the broad emission lines, producing an extended velocity component and artificially increasing SMBH masses by 1–2 orders of magnitude, when calculated using the virial SMBH estimator (Rusakov et al., 2025), although this scenario has been disputed for at least some JWST AGN (Juodžbalis et al., 2025). Overall, despite their unusual properties, it therefore appears likely that the majority of LRDs host AGN, potentially representing an early SMBH growth phase which would help explain their strong redshift dependence (Inayoshi, 2025) and overmassive SMBHs compared to their hosts; however, as for any broad-band photometric selection, it is quite likely there is some heterogeneity in the LRD population.
@@ 5.4.2 Galactic and large-scale environments of high-redshift quasars and JWST AGN
The majority of the AGN candidates detected by JWST were identified serendipitously in extensive spectroscopic surveys, in both average fields and quasar fields (with candidates in the foreground and background of the quasars). Despite the modest source statistics and cosmological volumes probed by the JWST surveys, it already seems likely that the JWST AGN reside in lower-mass haloes than the high-redshift quasars. Performing angular and projected cross-correlation analyses of 27 JWST AGN at with 679 photometrically selected galaxies, Arita et al. (2025) estimated a typical dark-matter halo mass for the JWST AGN of , an order of magnitude lower than that of faint and bright high-redshift quasars (Arita et al., 2023); see also §4.4 & 5.3. Two-point cross correlation analyses of LRDs with co-eval H emitting galaxies over 4–6 suggest that LRDs reside in similarly massive dark-matter halos of (1–2) (Lin et al., 2025); complementary constraints are also placed by Pizzati et al. (2024) who, on the basis of the source density of LRDs at 4–6 and the relative number of coeval dark-matter haloes, argued that LRDs must reside in haloes of to avoid a halo occupancy of %. These results indicate that the JWST AGN reside in dark-matter haloes an order of magnitude less massive than quasars (see §5.3), suggesting that they represent either an entirely different population or a much-earlier SMBH growth phase.
@@ 5.4.3 Origins of the seeds of massive black holes: models and observational constraints
Although most of the theoretical pathways for BH seed formation have been around for quite some time383838The paper of Martin Rees (1978) remains a solid and enlightening foundation of most contemporary theories (Rees, 1978), the last decade has been crucial to assess their detailed physical aspects (e.g., fragmentation of gas clouds or their collapse into protostars, accretion onto protostars and whether feedback processes inhibit growth) through cosmological simulations and semi-analytical models (e.g., Inayoshi and Tanaka, 2015; Chon et al., 2016; Regan et al., 2016; Wise et al., 2019; Latif et al., 2022).
@@
What have we learned from recent observations with JWST? So far, JWST has provided three key insights. First, it has pushed the redshift record for the earliest observed SMBHs in the Universe to 8.5–10.6; see Figs. 25 & 29. The SMBH masses of these systems are highly uncertain but broadly appear to lie in the range of – , measured from the broad emission lines for the 8.5 broad-line systems (see §5.1) and estimated from the AGN luminosity and assuming an Eddington ratio for the 10 narrow-line systems. However, these SMBH masses are broadly consistent with that required to explain the previously known high-redshift quasars (Schneider et al., 2023). In other words, in terms of their growth history, these systems at could be the progenitors of the high-redshift quasars (see discussion in Volonteri et al., 2023); see Fig. 31. The occupation fraction — the number of BHs in bins of galaxy properties such as stellar mass or velocity dispersion — can theoretically provide insights on the origin of BHs (Greene et al., 2020). However, more theoretical work is required to generate precise predictions for each potential BH formation channel, which can then be compared to observational data. Preliminary results from JWST suggest a potentially high number density of AGN (see §5.4.1), and as these systems and the host galaxies in which they reside become better understood, these observations will enable more accurate constraints on the AGN occupation fraction. One of the most striking discoveries by JWST is the potentially high mass ratios in a significant population of AGN across a wide redshift range (). If confirmed, these observations could suggest the presence of heavy seeds (Natarajan et al., 2024) and/or faster super Eddington or Eddington limited growth of SMBHs relative to the stellar populations of their host galaxies. However, the recent identification of a JWST broad-line AGN with in a metal-poor galaxy (% solar metallicity) argues against substantial super Eddington, or even Eddington limited accretion, for at least one system as the required gas inflow would drive significant star formation and hence metal enrichment (Maiolino et al., 2025). If a substantial fraction of the JWST AGN population are also found to reside in metal-poor galaxies then this would point towards more exotic BH formation scenarios such as primordial BHs to explain these systems (Maiolino et al., 2025). Therefore, at the time of writing, efficient growth of light BHs with phases of super-Eddington accretion, Eddington-limited growth of heavy seeds, and primordial origins of BHs, all remain viable options to explain the JWST AGN and it is quite likely that no one scenario explains the whole population (Schneider et al., 2023; Bennett et al., 2024; Dayal, 2024; Natarajan et al., 2024).
@@ Acknowledgements
We thank the referee for constructive and positive feedback and the following people for commenting on earlier drafts of this review and/or providing data and figures: Alastair Basden, Xiaohui Fan, Vicky Fawcett, Claire Greenwell, Dale Kocevski, Rohit Kondapally, Karen Leighly, James Matthews, John Paice, Krisztina Perger, Amy Rankine, Jan Scholtz, Matthew Temple, and Carolin Villforth. We extend a special thanks to Matthew Temple for producing Figure 27. DMA thanks the Science Technology Facilities Council (STFC) for support through the Durham consolidated grant (ST/T000244/1). JA, CMH, and LKM acknowledge funding from United Kingdom Research and Innovation (UKRI) grants MR/T020989/1, MR/V022830/1, and MR/Y020405/1, respectively. MH acknowledges funding from the Swiss SNSF Starting Grant (218032). RCH acknowledges support from NASA via grant numbers 80NSSC22K0862 and 80NSSC23K0485, and is grateful to the Durham Department of Physics and Center for Extragalactic Astronomy for hosting a very enjoyable and productive sabbatical visit. We thank the Matariki Network of Universities, of which both Dartmouth College and Durham University are partners, for financial support of our institutional meetings.

### v2

@@ 5.4.1 Increasing the demographics of high-redshift AGN
Many LRDs have a spectral break around the wavelength of the Balmer break (Labbe et al., 2024; Ma et al., 2024; Wang et al., 2024). Analysing a large sample of LRD JWST spectra, Setton et al. (2024) found the inflexion in the “V” shaped SED corresponds to the Balmer break in 50% of sources; see Fig. 26 for the break in the composite LRD spectrum. They argued this break is unlikely to occur for dust-reddened AGN with a range of dust extinctions and host-galaxy strengths and, adopting “Occam’s razor”, proposed the properties of LRDs are from a single origin, potentially due to late-type stars. However, a growing number of studies have argued that an AGN buried in an extremely dense cocoon of gas (what is sometimes referred to as a “black-hole star” scenario) would also produce such feature and could simultaneously explain the Balmer absorption lines seen in the spectra of some JWST AGN (e.g., Inayoshi and Maiolino, 2024; de Graaff et al., 2025; Ji et al., 2025; Naidu et al., 2025). A potential consequence of the large gas density is scattering of the broad emission lines, producing an extended velocity component and artificially increasing SMBH masses by 1–2 orders of magnitude, when calculated using the virial SMBH estimator (Rusakov et al., 2025), although this scenario has been disputed for at least some JWST AGN (Juodžbalis et al., 2025). Overall, despite their unusual properties, it therefore appears likely that the majority of LRDs host AGN, potentially representing an early SMBH growth phase which would help explain their strong redshift dependence (Inayoshi, 2025) and overmassive SMBHs compared to their hosts; however, as for any broad-band photometric selection, it is quite likely there is some heterogeneity in the LRD population (see, for example Setton et al., 2024; Zwick et al., 2025, for some alternative non-SMBH and non-AGN scenarios).
@@ 5.4.2 Galactic and large-scale environments of high-redshift quasars and JWST AGN
The majority of the AGN candidates detected by JWST were identified serendipitously in extensive spectroscopic surveys, in both average fields and quasar fields (with candidates in the foreground and background of the quasars). Despite the modest source statistics and cosmological volumes probed by the JWST surveys, it already seems likely that the JWST AGN reside in lower-mass haloes than the high-redshift quasars. Performing angular and projected cross-correlation analyses of 27 JWST AGN at with 679 photometrically selected galaxies, Arita et al. (2025) estimated a typical dark-matter halo mass for the JWST AGN of , an order of magnitude lower than that of faint and bright high-redshift quasars (Arita et al., 2023); see also §4.4 & 5.3. Two-point cross correlation analyses of LRDs with co-eval H emitting galaxies over 4–6 suggest that LRDs reside in similarly massive dark-matter halos of (1–2) (Lin et al., 2025); complementary constraints are also placed by Pizzati et al. (2024) who, on the basis of the source density of LRDs at 4–6 and the relative number of coeval dark-matter haloes, argued that LRDs must reside in haloes of to avoid a halo occupancy of %. These results indicate that the JWST AGN reside in dark-matter haloes an order of magnitude less massive than quasars (see §5.3), suggesting that they represent either an entirely different population or a much-earlier SMBH growth phase; however, we note that at least one 7 LRD is found in a much denser large-scale environment (Schindler et al., 2024), suggesting some diversity in LRD environments.
@@ 5.4.3 Origins of the seeds of massive black holes: models and observational constraints
Although most of the theoretical pathways for BH seed formation have been around for quite some time383838The paper of Martin Rees (1978) remains a solid and enlightening foundation of most contemporary theories (Rees, 1978)., the last decade has been crucial to assess their detailed physical aspects (e.g., fragmentation of gas clouds or their collapse into protostars, accretion onto protostars and whether feedback processes inhibit growth) through cosmological simulations and semi-analytical models (e.g., Inayoshi and Tanaka, 2015; Chon et al., 2016; Regan et al., 2016; Wise et al., 2019; Latif et al., 2022).
@@
What have we learned from recent observations with JWST? So far, JWST has provided three key insights. First, it has pushed the redshift record for the earliest observed SMBHs in the Universe to 8.5–10.6; see Figs. 25 & 29. The SMBH masses of these systems are highly uncertain but broadly appear to lie in the range of – , measured from the broad emission lines for the 8.5 broad-line systems (see §5.1) and estimated from the AGN luminosity and assuming an Eddington ratio for the 10 narrow-line systems. However, these SMBH masses are broadly consistent with that required to explain the previously known high-redshift quasars (Schneider et al., 2023). In other words, in terms of their growth history, these systems at could be the progenitors of the high-redshift quasars (see discussion in Volonteri et al., 2023); see Fig. 31. The occupation fraction — the number of BHs in bins of galaxy properties such as stellar mass or velocity dispersion — can theoretically provide insights on the origin of BHs (Greene et al., 2020). However, more theoretical work is required to generate precise predictions for each potential BH formation channel, which can then be compared to observational data. Preliminary results from JWST suggest a potentially high number density of AGN (see §5.4.1), and as these systems and the host galaxies in which they reside become better understood, these observations will enable more accurate constraints on the AGN occupation fraction. One of the most striking discoveries by JWST is the potentially high mass ratios in a significant population of AGN across a wide redshift range (). If confirmed, these observations could suggest the presence of heavy seeds (Natarajan et al., 2024) and/or faster super Eddington or Eddington limited growth of SMBHs relative to the stellar populations of their host galaxies. However, the recent identification of a JWST broad-line AGN with in a metal-poor galaxy (% solar metallicity) argues against substantial super Eddington, or even Eddington limited accretion, for at least one system as the required gas inflow would drive significant star formation and hence metal enrichment (Maiolino et al., 2025). If a substantial fraction of the JWST AGN population are also found to reside in metal-poor galaxies then this would point towards more exotic BH formation scenarios such as primordial BHs to explain these systems (Maiolino et al., 2025); see §5.4.1 for more details. Therefore, at the time of writing, efficient growth of light BHs with phases of super-Eddington accretion, Eddington-limited growth of heavy seeds, and primordial origins of BHs, all remain viable options to explain the JWST AGN and it is quite likely that no one scenario explains the whole population (Schneider et al., 2023; Bennett et al., 2024; Dayal, 2024; Natarajan et al., 2024).
@@ Acknowledgements
We thank the referee for constructive and positive feedback and the following people for commenting on earlier drafts of this review and/or providing data and figures: Alastair Basden, Xiaohui Fan, Vicky Fawcett, Claire Greenwell, Stephanie Juneau, Dale Kocevski, Rohit Kondapally, Karen Leighly, Rusen Lu, James Matthews, John Paice, Krisztina Perger, Cristina Ramos-Almeida, Amy Rankine, Jan-Torge Schindler, Jan Scholtz, Matthew Temple, Carolin Villforth, and Sarah White. We extend a special thanks to Matthew Temple for producing Figure 27. DMA thanks the Science Technology Facilities Council (STFC) for support through the Durham consolidated grant (ST/T000244/1). JA, CMH, and LKM acknowledge funding from United Kingdom Research and Innovation (UKRI) grants MR/T020989/1, MR/V022830/1, and MR/Y020405/1, respectively. MH acknowledges funding from the Swiss SNSF Starting Grant (218032). RCH acknowledges support from NASA via grant numbers 80NSSC22K0862 and 80NSSC23K0485, and is grateful to the Durham Department of Physics and Center for Extragalactic Astronomy for hosting a very enjoyable and productive sabbatical visit. We thank the Matariki Network of Universities, of which both Dartmouth College and Durham University are partners, for financial support of our institutional meetings.